# A New Method to Simulate Dark Matter–Baryon Interactions and Application to an Isolated Disk Galaxy

### 25 Mar 2026 v1

@@ IV.1 Impacts on morphology
The largest cross sections (300, 1000 mb) show that a considerable amount of gas falls into the center, forming an extremely compact and dense central disk of size $\sim 50\ {\rm pc}$. Because this size is approximately the minimum allowed softening length for gas particles and because we use a Jeans pressure floor, higher resolution is required to resolve the structure in the center. For 1000 mb, this process starts sooner and proceeds more quickly, eventually leading to the development of a small ring around the center with a radius of a few hundred parsec. The stellar disk also concentrates in the innermost $\sim 1\text{--}2\ {\rm kpc}$; because stars are collisionless, their response demonstrates that the potential changes drastically. Smaller cross sections show no significant differences from CDM.
Figure 2: Dark matter density profile at the end of the simulations (at about 1 Gyr). Plotted for reference are an NFW profile (dashed grey line) and a profile proportional to $r^{-2}$ (dotted grey line). The largest two cross sections (300 and 1000 mb) are much denser in the center than all other simulations, with 1000 mb becoming steeper even than $r^{-2}$.
@@
Figure 3: Enclosed dark matter (dashed) and gas (solid) mass profiles at the end (about 1 Gyr) of the simulations. Gas appears to concentrate in the center, with density increasing as a function of cross section. At the largest cross sections (300, 1000 mb), dark matter concentrates in the center as well.
@@ IV.2 Impacts on gas and stars
Figure 4: The median gas temperature computed in 1 kpc bins of cylindrical radius. For a cross section of 1000 mb, there is a significant temperature increase at all radii, but the effect is largely gone for cross sections below 300 mb.
Another quantity of interest is the temperature profile of the gas disk. For the final snapshot of each simulation, we compute the median gas temperature in 1 kpc bins of cylindrical radius up to $R=15\ {\rm kpc}$; the temperature profile is shown in Figure 4. We see no significant differences at small cross sections. With the 1000 mb simulation, however, the gas temperature is increased by a significant amount at most radii. We note that this is expected: in a Milky Way-like galaxy, the dark matter velocity dispersion is of order a few hundred km/s, which translates to an effective temperature of $\mathcal{O}(10^{6}\,\text{K})$ for $m_{\chi}=2\,m_{p}$. Hence, the dark matter is “hotter” than the gas and therefore heats it for sufficiently strong interactions.
@@
Because dark matter–baryon interactions both heat the gas and change its density, the effects on the star formation history of the galaxy may be non-trivial. To investigate that, for each snapshot we track the total mass in newly-formed stars, and take $\Delta M_{\star}/\Delta t$ as an estimate of the star formation rate. Figure 5 shows the evolution of the star formation rate over the history of the simulation. With 1000 mb, we see enhanced star formation rates starting almost immediately, despite the elevated gas temperature. At 300 mb, the star formation rate shows few differences from CDM until around 600 Myr, at which there is a short burst of increased star formation. At smaller cross sections, there are only small deviations from CDM which are uncorrelated with cross section, and so are likely just stochastic noise.
@@ IV.3 Possible role of bar formation
where $M_{j}$ is the mass of star particle $j$ and $\varphi_{j}$ is its azimuthal angle (Beane et al., 2023). For every snapshot of the simulations, we compute the bar strength $b$ in 1 kpc bins of cylindrical radius up to $R=15\ {\rm kpc}$ and record the maximum over the bins. Figure 6 shows the bar strength as a function of time, revealing that it evolves almost identically across all cross sections, reaching values above 0.2—a strong indicator for bar formation—within 300 Myr. In particular, during the period of bar formation (the first 300 Myr), we find no dependence of either the bar strength or speed of bar formation on cross section for these initial conditions.
@@
Having convinced ourselves that formation of the bar is not affected by the presence or absence of dark matter–baryon interactions, we next isolate the dark matter–baryon effects from bar-driven inflows. To do this, we perform a set of simulations with the stellar disk frozen in place; since the stars do not move, they cannot form a bar. The enclosed mass profiles of the end states of these simulations are shown in Figure 7. The results are very similar to the simulations without bar formation, indicating that dark matter–baryon interactions, and not bar formation, are driving gas and dark matter toward the center in this simulation. This does not mean that dark matter–baryon interactions have no effect on bar formation, and we believe this is an interesting question which merits future study.
@@ IV.4 Angular momentum
Figure 8: Evolution of the $z$-component of the total angular momentum for particles in the innermost 15 kpc of the galaxy. “Old stars” denote the stellar disk and bulge which are part of the initial conditions; “young stars” denote those which are formed out of the gas during the simulation. At large cross sections, the dark matter gains angular momentum relative to the CDM simulation, while the gas and young stars lose angular momentum. No significant differences can be seen between CDM and the smaller cross sections.
The angular momentum of the gas disk (in particular, its $z$-component) may play a role in the development of gas inflows. Dark matter–baryon interactions introduce a new possible mechanism for transferring angular momentum into or out of the gas disk. To explore this, we investigate the evolution of the $z$-component of the angular momentum during these simulations. Because the dark matter at large radii in general dominates the total angular momentum, we focus on the particles in the innermost 15 kpc. The evolution of the total angular momentum of each component in the innermost 15 kpc is shown in Figure 8. We see that at large cross section (1000 mb) the dark matter gains angular momentum relative to the CDM simulation, while the gas and newly-formed stars lose angular momentum.
@@
In a galaxy simulation, it is difficult to test for conservation of energy, because physics modules like radiative cooling allow for the dissipation of energy over time. However, angular momentum is conserved, which gives us a useful diagnostic for whether our treatment of the dark matter–baryon interactions has introduced a source of non-conservation. For each snapshot of our simulations, we compute the total angular momentum $\bm{L}=\sum_{i}M_{i}\,\bm{r}_{i}\times\bm{v}_{i}$ and evaluate how the magnitude of this vector evolves. The evolution is shown in Figure 9, revealing that there is a non-zero but small accumulation of angular momentum during the simulation across cross sections: about 0.03% per Gyr, even with dark matter–baryon interactions disabled, or about double that for the 1000 mb case. Some small angular momentum non-conservation may be expected due to the finite separation between macroparticles at the time of interactions (as is true for SIDM implementations, see Tulin and Yu, 2018), which would naturally be greater with large cross sections due to the increased number of scatterings. However, since there is some accumulation in the CDM simulation, it is not clear how much of the increase at 1000 mb should be attributed to our implementation versus the evolution of the strongly reshaped galaxy, which has far more extreme dynamics.
@@ V Summary and Conclusions
We have developed a method for incorporating dark matter–baryon interactions in cosmological $N$-body simulations which is accurate even in the challenging domain of dark matter with comparable or lower mass than the baryon. Our method treats phenomenological power-law cross sections of the form $\sigma=\sigma_{0}\,(v/c)^{n}$. We have implemented the method in the simulation code GIZMO (Hopkins, 2015). Our implementation is open-source, released under the GNU General Public License, and available on GitHub. [^5] https://github.com/cmhainje/gizmo-public We tested the accuracy of the model in a range of environments, including one test which is reproduced from Fischer et al. (2025) and reveals nearly identical behavior.
@@
We have applied this implementation to the evolution of an isolated Milky Way-like disk galaxy, using initial conditions from the AGORA project (Kim et al., 2016) and simulating the system for 1 Gyr. We considered a dark matter particle of mass $2\,m_{p}$ with an isotropic, velocity-independent DM–baryon cross section of 0, 10, 30, 100, 300, and 1000 mb. For the cases of 300 and 1000 mb, the DM–b interaction makes a clear change to the mass distribution in the central kiloparsec of the galaxy, significantly increasing the concentration of baryons and dark matter at the galaxy center in less than 1 Gyr. This occurs both with and without bar formation. Interactions at this strength also heat the gas, increase the star formation rate, and facilitate the transfer of angular momentum out of the gas disk.
@@
In our simulations with cross sections of 100 mb and less, the initial gas supply is largely used up before the DM–b interactions have any significant impact. This is not surprising since we do not simulate gas accretion into the galaxy. Properly simulating the effect of a DM–b cross section of order 10 mb or less—the range relevant for sexaquark dark matter, for instance—demands the use of high-quality hydrodynamics and detailed study of the fidelity of the subgrid models of stellar feedback. These efforts are not justified in this exploratory investigation of an idealized, isolated galaxy. In future work, we plan to study cosmological galaxy formation in the interesting cross section range, with the required sophisticated treatment of the “gastrophysics”.
@@
The code and method we have developed and presented here to incorporate dark matter–baryon interactions gives a practical, efficient, and reliable tool for the community. It should enable robust limits to be placed on various scenarios for dark matter–baryon interactions based on properties of galaxies.
@@
We thank Michael Blanton, Shy Genel, Chris Hayward, David W. Hogg, Clark Miyamoto, and Ben Wandelt for helpful discussions. We especially thank Angus Beane for his advice and debugging aid, and Phil Hopkins for illuminating conversations about GIZMO and the possibilities for implementing this new feature. CH is supported by the National Science Foundation Graduate Research Fellowship under Grant Number DGE-2234660. This research of GRF was supported in part by the Simons Foundation. This work made use of New York University IT High Performance Computing resources, services, and staff expertise.
@@ Appendix A Derivation of dark matter evolution
$f_{\chi}^{i}\bigg|_{t+\Delta t}=f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t$ (A3)
@@
The result contains two integrals over ${\rm d}^{3}v_{B}$ for each gas particle $j$. The second is tricky, as the integration variables $\bm{v}_{B}$ and $\Omega$ are used in the scattering transformation $\bm{S}_{\chi}$ inside the Dirac delta function. As such, it is expedient to approximate the integrals instead by Monte Carlo sampling. The integrals can be performed independently for each particle $j$ using single samples from the baryon Maxwell-Boltzmann distribution, which we denote $\bm{v}^{B}_{ij}$, and the angular component of the differential cross section, $\Omega_{ij}$. The result is
@@ Appendix B Derivation of baryon evolution
Because the baryons are assumed to thermalize on timescales faster than the timesteps of the simulation, we assume that the baryon velocity distribution is at all times Maxwell-Boltzmann. Thus, the evolution of gas particle $j$ is described by the rates-of-change of the bulk velocity and temperature:
@@
As we proceed, we will use the notation $f_{\chi}=f_{\chi}(\bm{x},\bm{v}_{\chi})$, $f_{\chi}^{\prime}=f_{\chi}(\bm{x},\bm{v}_{\chi}^{\prime})$, $f_{B}=f_{B}^{j}(\bm{x},\bm{v}_{B})$, $f_{B}^{\prime}=f_{B}^{j}(\bm{x},\bm{v}_{B}^{\prime})$.
We start with the acceleration. Inserting the Boltzmann equation gives
@@
$n_{j}(\bm{x})\,\frac{3}{\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}^{\prime}\,{\rm d}^{3}v_{B}^{\prime}\,{\rm d}\Omega\,f_{\chi}^{\prime}\,f_{B}^{\prime}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B^{\prime}\chi^{\prime}}\,\delta\big(\bm{v}_{B}-\bm{S}(\bm{v}_{B}^{\prime})\big)\,(\bm{v}_{B}-\bm{V}_{j})^{2}$
$\qquad-\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,(\bm{v}_{B}-\bm{V}_{j})^{2}$ (B11)
@@
$=\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,\big[(\bm{S}(\bm{v}_{B})-\bm{v}_{B})^{2}+2\,(\bm{v}_{B}-\bm{V}_{j})\cdot(\bm{S}(\bm{v}_{B})-\bm{v}_{B})\big]$ (B13)
@@
As noted by Ali-Haïmoud (2019), $\mathcal{A}$ and $\mathcal{B}$ have simple forms in terms of special functions if the momentum transfer cross section has a power law dependence on $v$. Specifically, if $\overline{\sigma}(v_{\rm rel})=\sigma_{0}\,(v_{\rm rel}/c)^{n}$ for integer $n\geq-4$, then

### 17 Apr 2026 v2

@@ IV.1 Impacts on morphology
The largest cross sections (300, 1000 mb) show that a considerable amount of gas falls into the center, forming an extremely compact and dense central disk of size $\sim 50\ {\rm pc}$. Because this size is approximately the minimum allowed softening length for gas particles and because we use a Jeans pressure floor, higher resolution is required to resolve the structure in the center. For 1000 mb, this process starts sooner and proceeds more quickly, eventually leading to the development of a small ring around the center with a radius of a few hundred pc. The stellar disk also concentrates in the innermost $\sim 1\text{--}2\ {\rm kpc}$; because stars are collisionless, their response demonstrates that the potential changes drastically. Smaller cross sections show no significant differences from CDM.
Figure 3: Dark matter density profile at the end of the simulations (at about 1 Gyr). Plotted for reference are an NFW profile chosen to match the initial conditions (dashed grey line) and a profile proportional to $r^{-2}$ (dotted grey line). The largest two cross sections (300 and 1000 mb) are much denser in the center than all other simulations, with 1000 mb becoming steeper even than $r^{-2}$.
@@
Figure 4: Enclosed dark matter (dashed) and gas (solid) mass profiles at the end of the simulations (at about 1 Gyr). Gas appears to concentrate in the center, with density increasing as a function of cross section. At the largest cross sections (300, 1000 mb), dark matter concentrates in the center as well.
@@ IV.2 Impacts on gas and stars
Figure 5: The median gas temperature at the end of the simulations, computed in 1 kpc bins of cylindrical radius. For a cross section of 1000 mb, there is a significant temperature increase at all radii, but the effect is largely gone for cross sections below 300 mb.
Another quantity of interest is the temperature profile of the gas disk. For the final snapshot of each simulation, we compute the median gas temperature in 1 kpc bins of cylindrical radius up to $R=15\ {\rm kpc}$; the temperature profile is shown in Figure 5. We see no significant differences at small cross sections. With the 1000 mb simulation, however, the gas temperature is increased by a significant amount at most radii. This is expected: in a Milky Way-like galaxy, the dark matter velocity dispersion is of order a few hundred km/s, which translates to an effective temperature of $\mathcal{O}(10^{6}\,\text{K})$ for $m_{\chi}=2\,m_{p}$. Hence, the dark matter is “hotter” than the gas and therefore heats it for sufficiently strong interactions.
@@
Because dark matter–baryon interactions both heat the gas and change its density, the effects on the star formation history of the galaxy may be non-trivial. To investigate that, for each snapshot we track the total mass in newly formed stars, and take the change in stellar mass divided by the time interval between snapshots as an estimate of the star formation rate. Figure 6 shows the evolution of the star formation rate over the history of the simulation. With 1000 mb, we see enhanced star formation rates starting almost immediately, despite the elevated gas temperature. At 300 mb, the star formation rate shows few differences from CDM until around 600 Myr, at which point there is a burst of increased star formation. At smaller cross sections, there are only small deviations from CDM that are uncorrelated with cross section, and so are likely just stochastic noise.
@@ IV.3 Possible role of bar formation
where $M_{j}$ is the mass of star particle $j$ and $\varphi_{j}$ is its azimuthal angle (Beane et al., 2023). For every snapshot of the simulations, we compute the bar strength $b$ in 1 kpc bins of cylindrical radius up to $R=15\ {\rm kpc}$ and record the maximum over the bins. Figure 7 shows the bar strength as a function of time, revealing that it evolves almost identically across all cross sections, reaching values above 0.2—a strong indicator for bar formation—within 300 Myr. In particular, during this initial period of bar formation, we find no dependence of either the bar strength or the speed of bar formation on cross section for these initial conditions.
@@
Having convinced ourselves that formation of the bar is not affected by the presence or absence of dark matter–baryon interactions, we next isolate the dark matter–baryon effects from bar-driven inflows. To do this, we perform a set of simulations with the stellar disk frozen in place; since the stars do not move, they cannot form a bar. The enclosed mass profiles of the end states of these simulations are shown in Figure 8. The results are very similar to the simulations with bar formation, indicating that dark matter–baryon interactions, and not bar formation, are driving gas and dark matter toward the center in this simulation. We emphasize that this does not mean that dark matter–baryon interactions have no effect on bar formation, and we consider this an interesting question that merits future study.
@@ IV.4 Angular momentum
Figure 9: Evolution of the $z$-component of the total angular momentum for particles in the innermost 15 kpc of the galaxy. “Old stars” denote the stellar disk and bulge that are part of the initial conditions; “young stars” denote those that are formed out of the gas during the simulation. At large cross sections, the dark matter gains angular momentum relative to the CDM simulation, while the gas and young stars lose angular momentum. No significant differences are seen between CDM and the smaller cross sections.
The angular momentum of the gas disk (in particular, its $z$-component) may play a role in the development of gas inflows. Dark matter–baryon interactions introduce a new possible mechanism for transferring angular momentum into or out of the gas disk. To explore this, we investigate the evolution of the $z$-component of the angular momentum during these simulations. Because the dark matter at large radii in general dominates the total angular momentum, we focus on the particles in the innermost 15 kpc. The evolution of the total angular momentum of each component in this region is shown in Figure 9. We see that at large cross section (1000 mb) the dark matter gains angular momentum relative to the CDM simulation, while the gas and newly formed stars lose angular momentum, providing evidence that DM–b interactions are facilitating a transfer of angular momentum.
@@
In a galaxy simulation, one cannot directly test for conservation of energy when using dissipative physics like radiative cooling. However, angular momentum is conserved up to errors from the tree gravity solver, which gives us a useful diagnostic for whether our treatment of the dark matter–baryon interactions has introduced a source of non-conservation. For each snapshot of our simulations, we compute the total angular momentum $\bm{L}=\sum_{i}M_{i}\,\bm{r}_{i}\times\bm{v}_{i}$ and evaluate how the magnitude of this vector evolves. The evolution is shown in Figure 10, revealing that there is a non-zero but small accumulation of angular momentum during the simulation across cross sections: about 0.03% per Gyr, even with dark matter–baryon interactions disabled, or about double that for the 1000 mb case. Some small angular momentum non-conservation may be expected due to the finite separation between macroparticles at the time of interactions (as is true for SIDM implementations, see Tulin and Yu, 2018), which would naturally be greater with large cross sections due to the increased number of scatterings. However, since there is some accumulation in the CDM simulation, it is not clear how much of the increase at 1000 mb should be attributed to our implementation versus the errors of the tree gravity solver in the strongly reshaped galaxy, which has a more extreme morphology.
@@ V Summary and conclusions
We have developed a method for incorporating dark matter–baryon interactions in cosmological $N$-body simulations that is accurate even in the challenging domain of dark matter with mass comparable to or lower than that of the baryon. We have implemented the method for isotropic power-law cross sections of the form $\sigma=\sigma_{0}\,(v/c)^{n}$, in the simulation code GIZMO (Hopkins, 2015). Our implementation is open-source, released under the GNU General Public License, and available on GitHub. [^9] https://github.com/cmhainje/gizmo-public We tested the accuracy of the model in a range of environments, including one test that is reproduced from Fischer et al. (2025) where it gives virtually identical behavior.
@@
We have applied our implementation to the evolution of an isolated Milky Way-like disk galaxy, using initial conditions from the AGORA project (Kim et al., 2016) and simulating the system for 1 Gyr. We considered a dark matter particle of mass $2\,m_{p}$ with isotropic, velocity-independent DM–b cross sections of 0, 10, 30, 100, 300, and 1000 mb. For the cases of 300 and 1000 mb, the DM–b interactions clearly change the mass distribution in the central kiloparsec of the galaxy, significantly increasing the concentration of baryons and dark matter. This occurs both with and without bar formation. Interactions at this strength also heat the gas, increase the star formation rate, and facilitate the transfer of angular momentum out of the gas disk.
@@
In our simulations with cross sections of 100 mb and less, the DM–b interactions do not have enough time to have any significant impact. This is not surprising, since the initial gas supply is partially used up, and we do not simulate gas accretion into the galaxy. Properly simulating the effect of a DM–b cross section of order 10 mb or less demands the use of high-quality hydrodynamics, modeling gas accretion, and the detailed study of the fidelity of the subgrid models of stellar feedback. These efforts are not justified in this exploratory investigation of an idealized, isolated galaxy. In future work, we plan to study cosmological galaxy formation in the interesting cross section range, with the required sophisticated treatment of the “gastrophysics”.
@@
The code and method we have developed and presented here to incorporate dark matter–baryon interactions provide a practical, efficient, and reliable tool for the community. It should enable robust limits to be placed on various scenarios for dark matter–baryon interactions based on properties of galaxies.
@@
We thank Michael Blanton, Shy Genel, Chris Hayward, David W. Hogg, Clark Miyamoto, and Ben Wandelt for helpful discussions. We thank Moritz Fischer for his helpful feedback on our discussion of the method he and his co-authors developed. We especially thank Angus Beane for his advice and debugging aid, and Phil Hopkins for illuminating conversations about GIZMO and the possibilities for implementing this new feature. CH is supported by the National Science Foundation Graduate Research Fellowship under Grant Number DGE-2234660. This research of GRF was supported in part by the Simons Foundation. This work made use of New York University IT High Performance Computing resources, services, and staff expertise.
@@ Appendix A Derivation of dark matter evolution
$f_{\chi}^{i}\bigg|_{t+\Delta t}=f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t.$ (A3)
@@
The result contains two integrals over ${\rm d}^{3}v_{B}$ for each gas particle $j$. The second is tricky, as the integration variables $\bm{v}_{B}$ and $\Omega$ are used in the scattering transformation $\bm{S}_{\chi}$ inside the Dirac delta function. Therefore, it is expedient to approximate the integrals instead by Monte Carlo sampling. The integrals can be performed independently for each particle $j$ using single samples from the baryon Maxwell-Boltzmann distribution, which we denote $\bm{v}^{B}_{ij}$, and the angular component of the differential cross section, $\Omega_{ij}$. The result is
@@ Appendix B Derivation of baryon evolution
Because the baryons are assumed to thermalize on timescales faster than the timesteps of the simulation, we impose that the baryon velocity distribution is at all times Maxwell-Boltzmann. Thus, the evolution of gas particle $j$ is described by the rates-of-change of the bulk velocity and temperature:
@@
As we proceed, we will use the notation $f_{\chi}=f_{\chi}(\bm{x},\bm{v}_{\chi})$, $f_{\chi}^{\prime}=f_{\chi}(\bm{x},\bm{v}_{\chi}^{\prime})$, $f_{B}=f_{B}^{j}(\bm{x},\bm{v}_{B})$, $f_{B}^{\prime}=f_{B}^{j}(\bm{x},\bm{v}_{B}^{\prime})$. We start with the acceleration. Inserting the Boltzmann equation gives
@@
$n_{j}(\bm{x})\,\frac{3}{\mu_{j}}\,\frac{{\rm d}T_{j}}{{\rm d}t}$ $=\begin{aligned} &\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}^{\prime}\,{\rm d}^{3}v_{B}^{\prime}\,{\rm d}\Omega\,f_{\chi}^{\prime}\,f_{B}^{\prime}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B^{\prime}\chi^{\prime}}\,\delta\big(\bm{v}_{B}-\bm{S}(\bm{v}_{B}^{\prime})\big)\,(\bm{v}_{B}-\bm{V}_{j})^{2}\\ &\qquad-\int{\rm d}^{3}v_{B}\,{\rm d}^{3}v_{\chi}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,(\bm{v}_{B}-\bm{V}_{j})^{2}\end{aligned}$ (B11)
@@
$=\int{\rm d}^{3}v_{\chi}\,{\rm d}^{3}v_{B}\,{\rm d}\Omega\,f_{\chi}\,f_{B}\,\frac{{\rm d}\sigma}{{\rm d}\Omega}\,v_{B\chi}\,\big[(\bm{S}(\bm{v}_{B})-\bm{v}_{B})^{2}+2\,(\bm{v}_{B}-\bm{V}_{j})\cdot(\bm{S}(\bm{v}_{B})-\bm{v}_{B})\big].$ (B13)
@@
As noted by Ali-Haïmoud (2019), $\mathcal{A}$ and $\mathcal{B}$ have simple forms in terms of special functions if the momentum transfer cross section has a power-law dependence on $v$. Specifically, if $\overline{\sigma}(v_{\rm rel})=\sigma_{0}\,(v_{\rm rel}/c)^{n}$ for integer $n\geq-4$, then